I know the following problem is supposed to lead to an eqaudratic equation, but I can't seem to get there. Please held !
Thanks!
4/3x-1/2(x+1)

Answer by sdmmadam@yahoo.com(530)

I know the following problem is supposed to lead to an eqaudratic equation, but I can't seem to get there. Please held !
Thanks!
4/3x-1/2(x+1)
The given problem does not seem to be a quadratic equation,and since the word equation is mentioned,probably the given problem is
4/(3x)-1/2(x+1) =0 ----(1)
Multiplying through out by 6x(x+1) which is the lcm of (3x) and 2(x+1)
4X2(x+1)-1X(3x) = 0 (zero mulitplied by anything is zero)
8(x+1)-3x=0 ----(2)
8x+8-3x= 0
(8x-3x)+8 = 0
5x+8 = 0
5x = -8
x = (-8)/5
Answer: x = (-8/5)
Verification: Putting x = (-8/5)
LHS = 8X[(-8/5)+1]-3X(-8/5) = 8X(-3/5)+24/5 = -24/5+24/5 =0 =RHS